Monday September 8
Tuesday September 9
Vacuum techs began letting the vacuum come up to normal atmosphere at 10 AM in certain areas of the Tevatron (TeV).
Operations lost vacuum readbacks for one area of the Recycler at 10:30 AM. In other Recycler areas, techs let the vacuum up to normal atmosphere.
Operations lost MI vacuum readbacks for the MI3 area and the MI52 septa control at 12:30 PM. Techs resolved a power problem at MI20 that took care of these and other problems.
Wednesday September 10
At 12:30 AM, operators responded to a fire alarm in the MI60 service building. With firemen escorts, operators investigated but found no sign of smoke or fire. The Fire Department disabled the sensor and reported it for repair.
Comments and Suggestions
|last modified 9/10/2003 email Fermilab|